OpenStudy (anonymous):

Convert to a fraction in lowest terms. 12.4%

OpenStudy (anonymous):

12.4 = 62/5

OpenStudy (anonymous):

\[12.4\%=\frac{1240}{100}=\frac{62}{5} \]